Aries:

What is the diameter of a circle with a circumference of 84.78 inches? Use 3.14 for π.

Arizona:

26.99

Aries:

okay

KyledaGreat:

The formula for the circumference of a circle is C = 2πr, where C is the circumference, π is the mathematical constant pi, and r is the radius of the circle. We can rearrange this formula to solve for the radius: r = C / (2π) Substituting the given circumference of 84.78 inches and using 3.14 for π, we get: r = 84.78 / (2 x 3.14) ≈ 13.5 Therefore, the radius of the circle is approximately 13.5 inches. To find the diameter, we simply multiply the radius by 2: d = 2 x r = 2 x 13.5 = 27 Therefore, the diameter of the circle with a circumference of 84.78 inches is approximately 27 inches.

Aries:

Can u help with another one plz?

Aries:

A mountain bike tire completes 20 revolutions and travels 136 feet. Rounded to the nearest inch, what is the diameter of the bike’s tire? Use 3.14 for π.

KyledaGreat:

We can use the formula for the circumference of a circle to solve this problem. The circumference is equal to the distance traveled by the tire during one revolution. Let d be the diameter of the tire, then the circumference is: C = πd If the tire completes 20 revolutions and travels 136 feet, then the total distance traveled by the tire is: total distance = 20 revolutions x circumference = 20πd We can now solve for the diameter d: 20πd = 136 d = 136 / (20π) d ≈ 1.09 feet To convert this to inches, we multiply by 12, since there are 12 inches in a foot: d ≈ 13.08 inches Rounding to the nearest inch, we get: d ≈ 13 inches Therefore, the diameter of the mountain bike tire is approximately 13 inches.